OpenStudy (meilendurcer):

3/4 divide 4/5= in simpilest form

OpenStudy (stamp):

3/4 div 4/5 = 3/4 mult 5/4

OpenStudy (meilendurcer):

\[\frac{ 3 }{ 4 } \div \frac{ 4 }{ 5 }\]

OpenStudy (stamp):

I understand. \[3/4*5/4\]

OpenStudy (meilendurcer):

i still don't get it

OpenStudy (dumbcow):

rule when dividing by a fraction is multiply by reciprocal so dividing by (1/2) is same as multiplying by 2

OpenStudy (meilendurcer):

online school